(π‘₯𝑦2+ 𝑦 βˆ’ π‘₯)𝑑π‘₯ + π‘₯(π‘₯𝑦 + 1)𝑑𝑦 = 0


1
Expert's answer
2022-05-31T08:47:10-0400
"\dfrac{\partial Q}{\partial x}=2xy+1=\dfrac{\partial P}{\partial y}"

The system of two differential equations that define the functionΒ "u(x, y)" is


"\begin{cases}
 \dfrac{\partial u}{\partial x}=xy^2+y-x \\ \\
 \dfrac{\partial u}{\partial y}=x^2y+x 
\end{cases}"

Integrate the first equation over the variableΒ "x"


"u=\int(xy^2+y-x)dx+\varphi(y)"

"=\dfrac{x^2y^2}{2}+xy-\dfrac{x^2}{2}+\varphi(y)"

Differentiate with respect to "y"


"\dfrac{\partial u}{\partial y}=x^2y+x+\varphi'(y) =x^2y+x"

"\varphi'(y) =0"

"\varphi(y)=-\dfrac{C}{2}"

Then


"u=\dfrac{x^2y^2}{2}+xy-\dfrac{x^2}{2}-\dfrac{C}{2}"

The general solution of the exact differential equation is given by


"x^2y^2+2xy-x^2=C"